The letter​ "t" makes up an estimated 10​% of a certain language. Assume this is still correct. A random sample of 1200 letters is taken from a randomly​ selected, large book and the​ t's are counted. Find the approximate probability that the random sample of 1200 letters will contain 8.8​% or fewer​ t's.

P( p̂ ≤ 0.088)=____________ (round to three decimal places as needed)
Answer #1

Solution

Given that,

\sigma\hat p =  \sqrt[p ( 1 - p ) / n] =  \sqrt [(0.1 * 0.9) / 1200 ] = 0.0087

P( \hat p \leq 0.088) =

= P[( \hat p - \mu \hat p ) / \sigma \hat p\leq (0.088 - 0.1) / 0.0087]

= P(z \leq -1.38)

= 0.084

P( \hat p \leq 0.088) = 0.084